CONSIDERATIONS FOR THE USE OF QUANTITATIVE FT-MIR REFERENCE LIBRARIES IN ENVI-RONMENTAL MEASUREMENTS

\underline{GRANT M. PLUMMER}, Rho Squared, PO Box 61536, Durham, NC 27715; THOMAS J. GEYER, Midwest Research Institute, 5520 Dillard Rd., Suite 1000, Cary, NC 27511; WILLIAM K. REAGEN, 3M Environmental Laboratory, 935 Bush Ave., Bldg. 2-3E-09, St. Paul, MN 55133.

Fourier transform mid-infrared spectroscopy has become a standard tool for performing quantitative analyses of environmental gas samples. Libraries of quantitative spectra for many compounds are now available from several sources. We offer some data and analytical results pertaining to the use of these libraries in practical applications.